# Decoupled-Value Attention for Prior-Data Fitted Networks: GP Inference for Physical Equations

### v1

@@ 4.2 Physics Equation Learning
Rosenbrock Function: To benchmark our method against a well–established baseline, we conduct experiments on the 5-dimensional Rosenbrock function Rosenbrock (1960), a standard test problem in optimization that is often interpreted as a nonlinear potential energy landscape with a curved valley structure Akian et al. (2022). GPs are a natural choice for such comparisons because they provide a flexible non-parametric model with uncertainty quantification, and they have been widely benchmarked on Rosenbrock and related test functions in the GP literature Xu et al. (2025). Results indicate that 5D PFN with Transformer+DVA shows MSE 6.86.8e-4 and CNN+DVA achieves MSE of 1.61.6e-3, without any retraining, see Table 8 in Appendix D for detailed results.
Power Flow Learning: In this experiment, we model the IEEE 33-bus distribution system by treating the real and reactive power demands at each of the 32 load buses as uncertain inputs (same experiment design as described in Pareek & Nguyen (2021); Liu & Srikantha (2022)). This results in a 64-dimensional input space (32 active + 32 reactive loads). Now the learning task is to predict the corresponding steady-state bus voltage magnitude— effectively learning the nonlinear AC power flow mapping from loads to voltages i.e. $\text{Voltage}=f(\text{Loads})$ (See equation 10 in Appendix A.2). Table 3 benchmarks power flow surrogates under varying load perturbations from 5% to 50%. Exact GP achieves the lowest MSE and MAE values across all cases, but requires training 32 times (one for each node), which becomes infeasible for repeated queries under changing load conditions. In contrast, both PFNs CNN+DVA and Transformer+DVA trade a modest increase in error for dramatic efficiency gains—over $80\times$ faster than GPs—while maintaining voltage prediction accuracy at the order of $10^{-3}$, sufficient for practical grid analysis. Further, the prediction error decreases as more training (context) samples are provided, with both CNN+DVA and Transformer+DVA converging to near-identical performance as illustrated in Figure 7 (Appendix D). These results highlight that while GPs remain the gold standard for accuracy, DVA-equipped PFNs offer a scalable alternative, enabling high-dimensional, uncertainty-aware power flow learning in real time for complex networked systems. Moreover, because voltages are in per-unit, MSE and MAE values around $10^{-3}$ are practically acceptable. In real systems, measurement devices typically have least counts of $10^{-3}$ p.u., so an error of $10^{-3}$ in a 1 kV system corresponds to just 1 V Molzahn et al. (2019). We also note that, consistent with the 10D case in Figure 2, PFNs equipped with vanilla attention failed to train sufficiently for this 64D problem and thus did not yield meaningful results. Training time for 64D models is approximately 14 hours for both Transformer and CNN-based PFNs on NVIDIA 4500ADA GPU.
@@ 5 Conclusions and Future Work
Future work will focus on scaling PFNs to even larger power networks and higher-dimensional uncertainty spaces, particularly in the context of power flow uncertainty quantification and planning problems. Another promising direction is the design of architectures that explicitly support input-space localization beyond attention mechanisms, allowing PFNs to capture neighborhood structure in physical systems more faithfully. We would also like to work on DVA limitation of having no output affinities, while maintaining localization. Together, these efforts can push PFNs toward practical deployment for real-time, uncertainty-aware decision making in modern power systems.
6 Acknowledgment
The authors thank Dr. Sidhant Misra (LANL), Dr. Kaarthik Sundar (LANL), Dr. Abhijith Jayakumar (LANL), and Dr. Deepjyoti Deka (MIT) for discussions on physical system learning and for endorsing the hypothesis that CNNs can also do Bayesian Inference.
@@ A.1 Rosenbrock Function
For our baseline experiments, we use the 5D–dimensional Rosenbrock function, defined as Rosenbrock (1960)
@@
We normalize the input vectors $\mathbf{x}$ to the unit hypercube $0,1]^{5}$ before training, and outputs are standardized using $Z$-score normalization. For GP testing, we employ a Gaussian process surrogate with a RBF kernel with automatic relevance determination (ARD) length-scales Williams & Rasmussen (2006).
@@ A.2 AC Power Flow Problem
In our setting, we explicitly consider uncertainty at each bus in both real and reactive power injections. For an IEEE 33-bus system Pareek & Nguyen (2021); Liu & Srikantha (2022), with the first bus designated as the slack bus (zero load), this leads to a $64$-dimensional uncertainty input vector capturing nodal variability across all other buses. We follow the standard ACPF model used in PowerModels Coffrin et al. (2018), and we use compute_ac_pf function of PowerModels.jl to generate dataset.
@@ A.2.1 Power Flow Learning with GPs
where $y(\bm{x})$ is the observed voltage at a node for load vector $\bm{x}$. With GP priors, $\bm{y}(\bm{x})\sim\mathcal{GP}(0,K(\bm{x},\bm{x})+\sigma^{2}_{\varepsilon}I)$, and the kernel $K$ encodes correlations between operating points. Owing to the smoothness of voltages as a function of load, the squared exponential kernel has been widely used in prior work Tan et al. (2025).
@@
GP-based approximations have been shown to outperform analytically approximated linearized models in capturing power flow uncertainty Pareek & Nguyen (2021) and are favored over other learning methods such as DNN due to closed-form approximation nature of GP, and predictive variance availability etc. For more details on power flow modeling and GP surrogate of it, readers can refer to Tan et al. (2025).
Appendix B Implementation and Architectural Details
@@
Synthetic Prior Data Generation: To assess different attention mechanisms in Prior-Data Fitted Networks (PFNs), we use synthetic datasets generated from GP priors, following Müller et al. (2022). Each regression dataset consists of input–output pairs $(x,y)$, with inputs sampled uniformly and outputs drawn from a multivariate Gaussian with an RBF kernel having lengthscale, variance (output scale), and observation noise variance as hyperparameters. The inputs are normalized using Z-score normalization, while outputs are shifted to a range of 0.80.81.21.2 for all datasets.
@@
For classification-based objectives, the continuous output space is discretized into buckets derived from quantiles of GP-sampled outputs Müller et al. (2022). Each bucket corresponds to a categorical class index, allowing regression-style PFN training to be cast into classification under a Riemannian distribution loss formulation. This strategy preserves ordering structure while making outputs compatible with categorical training setups. See Müller et al. (2022) for more information in this.
@@
Hyperparameters Selection: To ensure fair evaluation across architectures and embedding dimensions, we employed Optuna Akiba et al. (2019) for automated hyperparameter tuning. Key parameters such as model width, hidden dimension size, number of attention blocks, number of heads, and dropout rate were jointly optimized, with AdamW and a linear warmup followed by step-wise decay. Each trial involved computing training and validation losses on PFN tasks, and Optuna’s pruning strategies enabled efficient exploration of the search space. The best-performing configurations were selected based on initial validation loss over 1000 trials, while training loss was also tracked to assess stability.
@@
Parameter Calculation: To compute the total learnable parameters (Table 7), we loaded each model’s state dictionary, where keys represent layers and values are tensors of weights or biases. For each tensor, we used numel() from pytorch and summed these values to obtain the total parameter count.
Appendix C Kernel-based Attention Experiment Details
For the experiments shown in Figure 1, we generated synthetic datasets using both smooth and non-smooth kernels, and trained 1D Transformer and CNN models (with the same backbone architectures as used in Table 2) for both kernel attention and DVA. The smooth datasets were sampled from an RBF kernel, which promotes locality and smoothness in the function. In contrast, the non-smooth datasets were generated using a linear–periodic kernel as discussed in Duvenaud (n.d.), which combines a linear trend with a periodic component, producing oscillatory patterns with irregular variations and reduced smoothness.
@@ Appendix C Kernel-based Attention Experiment Details
Appendix D Additional Results
Figure 5: Comparison with GP for 1D & 2D PFNs: MSE as a function of context size. All models are tested using $n_{\text{test}}=500$, for $N_{context}$. The results show that error consistently decreases with larger context sizes, and that CNN- and Transformer-based PFNs with DVA approach the performance of exact GP inference even in higher dimensions. Exact GP baselines were fit using scikit-learn with $N_{context}$ training samples.

### v2

@@ 4.2 Physics Equation Learning
Rosenbrock Function: To benchmark our method against a well–established baseline, we conduct experiments on the 5-dimensional Rosenbrock function rosenbrock1960, a standard test problem in optimization that is often interpreted as a nonlinear potential energy landscape with a curved valley structure akian2022learning. GPs are a natural choice for such comparisons because they provide a flexible non-parametric model with uncertainty quantification, and they have been widely benchmarked on Rosenbrock and related test functions in the GP literature xu2025standard. Results indicate that 5D PFN with Transformer+DVA shows MSE 6.86.8e-4 and CNN+DVA achieves MSE of 1.61.6e-3, without any retraining, see Table 9 in Appendix D for detailed results.
Power Flow Learning: In this experiment, we model the IEEE 33-bus distribution system by treating the real and reactive power demands at each of the 32 load buses as uncertain inputs (same experiment design as described in pareek2021framework; 9734745). This results in a 64-dimensional input space (32 active + 32 reactive loads). Now the learning task is to predict the corresponding steady-state bus voltage magnitude— effectively learning the nonlinear AC power flow mapping from loads to voltages i.e. $\text{Voltage}=f(\text{Loads})$ (See equation 9 in Appendix A.2). Table 4 benchmarks power flow surrogates under varying load perturbations from 5% to 50%. Exact GP achieves the lowest MSE and MAE values across all cases, but requires training 32 times (one for each node), which becomes infeasible for repeated queries under changing load conditions. In contrast, both PFNs CNN+DVA and Transformer+DVA trade a modest increase in error for dramatic efficiency gains—over $80\times$ faster than GPs—while maintaining voltage prediction accuracy at the order of $10^{-3}$, sufficient for practical grid analysis. Further, the prediction error decreases as more training (context) samples are provided, with both CNN+DVA and Transformer+DVA converging to near-identical performance as illustrated in Figure 9 (Appendix D). These results highlight that while GPs remain the gold standard for accuracy, DVA-equipped PFNs offer a scalable alternative, enabling high-dimensional, uncertainty-aware power flow learning in real time for complex networked systems. Moreover, because voltages are in per-unit, MSE and MAE values around $10^{-3}$ are practically acceptable. In real systems, measurement devices typically have least counts of $10^{-3}$ p.u., so an error of $10^{-3}$ in a 1 kV system corresponds to just 1 V molzahnsurvey. We also note that, consistent with the 10D case in Figure 2, PFNs equipped with vanilla attention failed to train sufficiently for this 64D problem and thus did not yield meaningful results. Training time for 64D models is approximately 14 hours for both Transformer and CNN-based PFNs on NVIDIA 4500ADA GPU.
@@ A.1 Rosenbrock Function
For our baseline experiments, we use the 5D–dimensional Rosenbrock function, defined as rosenbrock1960
@@
We normalize the input vectors $\mathbf{x}$ to the unit hypercube $0,1]^{5}$ before training, and outputs are standardized using $Z$-score normalization. For GP testing, we employ a Gaussian process surrogate with a RBF kernel with automatic relevance determination (ARD) length-scales williams2006gaussian.
@@ A.2 AC Power Flow Problem
In our setting, we explicitly consider uncertainty at each bus in both real and reactive power injections. For an IEEE 33-bus system pareek2021framework; 9734745, with the first bus designated as the slack bus (zero load), this leads to a $64$-dimensional uncertainty input vector capturing nodal variability across all other buses. We follow the standard ACPF model used in PowerModels powermodels, and we use compute_ac_pf function of PowerModels.jl to generate dataset.
@@ A.2.1 Power Flow Learning with GPs
where $y(\bm{x})$ is the observed voltage at a node for load vector $\bm{x}$. With GP priors, $\bm{y}(\bm{x})\sim\mathcal{GP}(0,K(\bm{x},\bm{x})+\sigma^{2}_{\varepsilon}I)$, and the kernel $K$ encodes correlations between operating points. Owing to the smoothness of voltages as a function of load, the squared exponential kernel has been widely used in prior work tan2025gaussian.
@@
GP-based approximations have been shown to outperform analytically approximated linearized models in capturing power flow uncertainty pareek2021framework and are favored over other learning methods such as DNN due to closed-form approximation nature of GP, and predictive variance availability etc. For more details on power flow modeling and GP surrogate of it, readers can refer to tan2025gaussian.
Appendix B Proof of DVA Localization Theorem
This section provides the full proof of Theorem 1 and Theorem 2 and an additional corollary covering the linear-embedding case, where the DVA logit are shown to be a Mahalanobis RBF kernel.
Proof of Theorem 1. Assume $u(\mathbf{x})=W_{q}W_{x}\mathbf{x}$ and $v(\mathbf{x})=W_{k}W_{x}\mathbf{x}$, and define $A=(W_{q}W_{x})^{\top}(W_{k}W_{x})$. Consider $A$ is symmetric and positive definite (e.g. Consider a case where $W_{q}=W_{k}$ with full rank $W$’s), then for any context point $\mathbf{x}_{i}$,
$\ell_{i}\;=\;\langle Q_{\star},K_{i}\rangle\;=\;\mathbf{x}_{\star}^{\top}A\mathbf{x}_{i}=\frac{1}{2}\big(\|\mathbf{x}_{\star}\|_{A}^{2}+\|\mathbf{x}_{i}\|_{A}^{2}-\|\mathbf{x}_{\star}-\mathbf{x}_{i}\|_{A}^{2}\big),$
with $\|z\|_{A}^{2}=z^{\top}Az$ being energy norm. Exponentiating and dividing by temperature $\tau$ ($\sqrt{d_{k}}$ in our case) gives
$\exp\!\big(\ell_{i}/\tau\big)=\exp\!\Big(\frac{\|x_{\star}\|_{A}^{2}+\|x_{i}\|_{A}^{2}}{2\tau}\Big)\cdot\exp\!\Big(-\frac{\|x_{\star}-x_{i}\|_{A}^{2}}{2\tau}\Big).$
Up to per-point norm factors, the attention weight from equation 5 satisfies
$\alpha_{i}(\mathbf{x}_{\star})\;\propto\;\exp\!\left(-\tfrac{1}{2}\|\mathbf{x}_{\star}-\mathbf{x}_{i}\|_{A}^{2}\right).$
Hence, DVA reduces to a Mahalanobis RBF kernel and is therefore automatically local: the attention weights decay exponentially with distance. $\square$
Proof of Theorem 2. Fix a query $\mathbf{x}_{\star}$ and context inputs $\{\mathbf{x}_{i}\}_{i=1}^{N_{\texttt{context}}}$ Let, $u(x)=W_{q}\,\varphi_{x}(x),\quad v(x)=W_{k}\,\varphi_{x}(x)$. Consider the query-key inner product as:
$\ell_{i}\;=\;\langle u(\mathbf{x}_{\star}),v(\mathbf{x}_{i})\rangle,\qquad\alpha_{i}(\mathbf{x}_{\star})=\frac{\exp(\ell_{i})}{\sum_{j=1}^{N_{\texttt{context}}}\exp(\ell_{j})}.$
For a fixed $\varepsilon>0$, define the sets of near and far indices
$N_{\varepsilon}=\{i:\|\mathbf{x}_{i}-\mathbf{x}_{\star}\|\leq\varepsilon\},\qquad F_{\varepsilon}=\{i:\|\mathbf{x}_{i}-\mathbf{x}_{\star}\|>\varepsilon\}.$
By the local-separation of inner product assumption, there exist constants $\gamma\in\mathbb{R}$ and $\delta>0$ such that $\ell_{i}\geq\gamma\,(i\in N_{\varepsilon}),$ and $\ell_{j}\leq\gamma-\delta\,(j\in F_{\varepsilon}).$ Hence, considering each exponent at bound
$\sum_{j\in F_{\varepsilon}}\exp(\ell_{j})\;\leq\;|F_{\varepsilon}|\,e^{\gamma-\delta},\qquad\sum_{i\in N_{\varepsilon}}\exp(\ell_{i})\;\geq\;|N_{\varepsilon}|\,e^{\gamma}.$
Using these inequalities in the definition of the softmax weights yields
$\sum_{j\in F_{\varepsilon}}\alpha_{j}(\mathbf{x}_{\star})=\frac{\exp(\ell_{i})}{\sum_{i\in N_{\varepsilon}}\exp(\ell_{i})+\sum_{j\in F_{\varepsilon}}\exp(\ell_{j})}\;\leq\;\frac{\exp(\ell_{i})}{\sum_{i\in N_{\varepsilon}}\exp(\ell_{i})}=\frac{|F_{\varepsilon}|\,e^{\gamma-\delta}}{|N_{\varepsilon}|\,e^{\gamma}}=\frac{|F_{\varepsilon}|}{|N_{\varepsilon}|}\,e^{-\delta}.$
Therefore, as context length goes to infinity $N_{\texttt{context}}\rightarrow\infty$, local samples $|N_{\varepsilon}|\to\infty$ in probability while $|F_{\varepsilon}|/|N_{\varepsilon}|$ remains bounded ($F_{\varepsilon}$ is complement of $N_{\varepsilon}$). Therefore the right-hand side converges to 0 in probability, and thus
$\sum_{j\in F_{\varepsilon}}\alpha_{j}(\mathbf{x}_{\star})\;\xrightarrow[N_{\texttt{context}}\to\infty]{\mathbb{P}}\;0.$
This proves that DVA attention becomes fully input-local as context size grows. $\square$
Note[Local Separability Assumption]: The local-separability assumption is generally mild and naturally satisfied when the encoder $\varphi_{x}$ preserves the local geometry of the input space, i.e., when nearby inputs remain close in the embedding and distant inputs remain well separated. This typically holds for smooth, Lipschitz neural encoders such as MLPs or Fourier-feature mappings applied to regression tasks with locally regular structure. In contrast, the assumption may fail when the underlying task or encoder is inherently nonlocal—for example, when the target function is periodic, symmetric, or globally dependent, or when the encoder distorts geometry through aliasing or collapse. Simple cases include encoder collapse (e.g., when nonlinearities saturate and map many inputs to nearly the same vector) and extreme compression (e.g., a narrow bottleneck that forces very different inputs to share the same low-dimensional code). In these situations, distant inputs may appear artificially similar to the query, breaking the monotonic relationship between distance and attention.
Figure 4: In the first layer, we can clearly see that as the Euclidean distance increases, the softmax weight decreases exponentially, clearly showing that the DVA mechanism enforces localization. Localization is enforced in the first layer, and since layer 2 is the last layer of the model, which outputs the exact values, the softmax values are all minimal and closer to each other, suggesting the last layer averages the result for proper PPD approximation. Essentially, Layer 1 performs the “Local Smoothing” (gathering information from neighbors). Layer 2 performs “Feature Mixing” (processing the gathered information). Since Layer 1 has already gathered the local information into the latent vector, Layer 2 no longer needs to be spatially local; it can attend globally or uniformly to refine the prediction. This provided empirical evidence of the DVA localization theorem.
Figure 5: In contrast to the DVA, the standard VA Transformer assigns near-uniform attention weights across the entire input domain (flat trend lines). This confirms that without explicit localization through architecture changes (which we made in DVA), the model defaults to global averaging rather than local interpolation. Attention weights in both layers remain effectively uniform regardless of input distance, empirically validating Theorem 6.3 of Naglar 2023.
Appendix C Implementation and Architectural Details
@@
Synthetic Prior Data Generation: To assess different attention mechanisms in Prior-Data Fitted Networks (PFNs), we use synthetic datasets generated from GP priors, following müller2024transformersbayesianinference. Each regression dataset consists of input–output pairs $(x,y)$, with inputs sampled uniformly and outputs drawn from a multivariate Gaussian with an RBF kernel having lengthscale, variance (output scale), and observation noise variance as hyperparameters. The inputs are normalized using Z-score normalization, while outputs are shifted to a range of 0.80.81.21.2 for all datasets.
@@
For classification-based objectives, the continuous output space is discretized into buckets derived from quantiles of GP-sampled outputs müller2024transformersbayesianinference. Each bucket corresponds to a categorical class index, allowing regression-style PFN training to be cast into classification under a Riemannian distribution loss formulation. This strategy preserves ordering structure while making outputs compatible with categorical training setups. See müller2024transformersbayesianinference for more information in this.
@@
Hyperparameters Selection: To ensure fair evaluation across architectures and embedding dimensions, we employed Optuna akiba2019optunanextgenerationhyperparameteroptimization for automated hyperparameter tuning. Key parameters such as model width, hidden dimension size, number of attention blocks, number of heads, and dropout rate were jointly optimized, with AdamW and a linear warmup followed by step-wise decay. Each trial involved computing training and validation losses on PFN tasks, and Optuna’s pruning strategies enabled efficient exploration of the search space. The best-performing configurations were selected based on initial validation loss over 1000 trials, while training loss was also tracked to assess stability.
@@
Appendix D Additional Results
Figure 7: Comparison with GP: MSE for 1D,2D, 5D and 10D PFNs as a function of context size. All models are tested using $n_{\text{test}}=500$, for $N_{context}$. The results show that error consistently decreases with larger context sizes, and that CNN- and Transformer-based PFNs with DVA approach the performance of exact GP inference even in higher dimensions. Exact GP baselines were fit using scikit-learn with $N_{context}$ training samples.
@@ Appendix D Additional Results
Appendix E Kernel-based Attention Experiment Details
For the experiments shown in Figure 1, we generated synthetic datasets using both smooth and non-smooth kernels, and trained 1D Transformer and CNN models (with the same backbone architectures as used in Table 2) for both kernel attention and DVA. The smooth datasets were sampled from an RBF kernel, which promotes locality and smoothness in the function. In contrast, the non-smooth datasets were generated using a linear–periodic kernel as discussed in duvenaud_kernel_cookbook, which combines a linear trend with a periodic component, producing oscillatory patterns with irregular variations and reduced smoothness. Our results confirm that while Kernel-based attention and the proposed DVA reach effectively the same performance level on RBF training priors, Kernel attention requires significantly more computational time to do so. Quantitatively, both methods achieve comparable predictive error on 1D smooth functions (MSE of $1.40\times 10^{-4}$ for Kernel vs. $1.90\times 10^{-4}$ for DVA) and converge to similar validation losses on the 10D task (1.50-1.50 for Kernel vs. 1.29-1.29 for DVA). However, the training logs highlight a drastic difference in efficiency: DVA is approximately $\mathbf{4\times}$ faster, completing 80,000 optimization steps in just 42 minutes, whereas the Kernel model managed only 23,000 steps in a longer duration of 48 minutes, demonstrating that DVA delivers the performance of kernel methods with vastly superior throughput.
Appendix F Ablation Studies
F.1 Ablation of the Value Encoder $\phi_{y}$
This experiment studies whether a learnable value encoder $\phi_{y}$ is required and whether higher-capacity encoders improve GP-style inference in DVA-PFNs. A simple linear encoder from $y$ to the model dimension performs the best, matching a 2-layer MLP while substantially outperforming a direct broadcast baseline. These results highlight that a learnable value encoder is essential for mitigating the mismatch between GP regression weights (which can be positive or negative) and softmax attention weights (strictly positive). The encoder $\phi_{y}$ provides the necessary representational flexibility so that the model can emulate signed GP-like contributions through feature transformations rather than through the attention weights alone. Thus, although simple, the linear $\phi_{y}$ is necessary for effective DVA aggregation.
F.2 Ablation of the Input Encoder $\phi_{x}$
We evaluate whether a learnable input encoder is essential for representing the geometry implied by the GP prior. Both the linear encoder and the 2-layer MLP encoder perform nearly identically (validation loss $\approx-3.12$), indicating that moderate capacity is enough for expressing the input geometry required by DVA. Removing the encoder and broadcasting $x$ across channels significantly degrades performance (validation loss $\approx-1.68$), demonstrating that DVA needs a trainable representation but not additional encoder depth.
Figure 11: Ablation results for encoders $\phi_{y}$, $\phi_{x}$, the prediction head $g$, and query/key projections $W_{q},W_{k}$. Linear encoders for both $x$ and $y$ perform as well as deeper MLP-based encoders, whereas direct broadcasting produces large degradation. Learnable $\phi_{y}$ and $g$ are essential for mitigating the mismatch between GP regression weights (which may be positive or negative) and softmax attention weights (strictly positive). Finally, tying $W_{q}=W_{k}$ performs nearly identically to learning them independently, indicating that DVA-PFNs are robust to the choice of kernel-projection parameterization. Here $N_{\text{train}}$ is measured in millions.
F.3 Ablation of the Prediction Head $g(\cdot)$
We ablate the prediction head $g(\cdot)$, which converts the final hidden representation into bucketized predictive logits. Both the MLP head and the simple linear head converge near 3.1-3.1, indicating that additional nonlinearity does not improve predictive performance. However, removing all learnable parameters and broadcasting a single hidden dimension to all output bins fails completely (validation loss $\approx-0.90$). A learnable $g$ is therefore required for the same reason as $\phi_{y}$: GP regression weights can be positive or negative, while softmax attention weights cannot. The output head compensates by reparameterizing hidden features so that the model can emulate signed GP-like effects through linear combinations of positive-weighted attention outputs. Thus, any learnable $g(\cdot)$ is sufficient, but some learnability is essential.
F.4 Ablation of Query/Key Projections $W_{q},W_{k}$
This ablation tests whether DVA depends on having independent query and key projections. We compare the standard configuration with separate $W_{q}$ and $W_{k}$ to a shared-projection variant where $Q=K=W_{qk}(x)$. Across all training-set sizes (with $N_{\text{train}}$ measured in millions), both variants perform nearly identically, converging near 3.0-3.0. These findings show that DVA-PFN performance is not sensitive to whether query and key projections are shared or independent. The model appears to operate primarily through the relative input geometry encoded in $\phi_{x}$, not through asymmetry introduced by $W_{q}$ and $W_{k}$. This suggests redundancy in the standard attention parameterization and further indicates robustness of DVA to architectural choices.
F.5 Bucket Size Ablation
For both CNN+DVA and Transformer+DVA, increasing the number of output bins improves validation loss in both 1D (left column) and 2D (right column) tasks. The improvement is sharp when moving from small to moderate bin sizes, and performance stabilizes once the bin size reaches roughly 150–200 bins. Overall, DVA-PFNs are not highly sensitive to bin size, as long as it is sufficiently large to provide adequate resolution of the predictive distribution.
Figure 12: Effect of Bin Size on Validation Loss: Validation loss variation with different bin sizes for CNN+DVA (top row) and Transformer+DVA (bottom row) across 1D (first column), and 2D (second column).
Appendix G Linear Attention
Figure 13: Comparison with performer/linear attention using Transformer PFN. Linear attention is used with both VA and DVA. As is visible, the decoupling is allowing for better training; the effect is visible clearly in 10D. Softmax DVA performs best as expected, followed by Linear approximation of DVA. VA performs worse, and is unable to train in 10D.
Appendix H Predictive Uncertainty Calibration Results
Figure 14: Comparing the PPD for three distinct samples for (Top) Bus 18. (Bottom) Bus 33
Table 10: Percentage of values within variance bounds node 32 (farthest node from Generator bus) and 4500 testing samples (500 Context/Training points)
Model $\pm 0.1\sigma$ $\pm 1\sigma$ $\pm 2\sigma$
GP 48.36% 99.93% 100.00%
Transformer + DVA 7.73% 67.11% 96.13%
CNN + DVA 12.07% 89.00% 99.91%
Appendix I Consolidated Hyperparameters
To ensure reproducibility, we provide the detailed hyperparameters used for synthetic data generation, model architectures (Transformer, CNN, RNN, LSTM), and the optimization process. All models were implemented in PyTorch.
Synthetic Data Generation (GP Prior)
For all experiments, training data was generated on-the-fly using Gaussian Process priors with Radial Basis Function (RBF) kernels. Output values $y$ were shifted to ensure positive support for the bucketization process. Table 15 lists the kernel parameters for each dimensional setting.
Model Architecture
Transformer models use standard Multi-Head Attention (or the proposed DVA modification) with Pre-LayerNorm. The architectural hyperparameters were tuned
CNN models map inputs to an embedding space and process them with 1D convolution blocks over the sequence dimension.
Table 11: Hyperparameters for Synthetic GP Prior Data Generation.
Input Dim Points / Dataset ($N$) Kernel Type Length-scale ($\ell$) Kernel Var. Noise ($\sigma^{2}$) Output Shift
1D 100 RBF 0.6 0.01 $1\times 10^{-2}$ 1.0
2D 100 RBF 0.6 0.01 $1\times 10^{-2}$ 1.0
5D 400 RBF 0.6 0.001 $1\times 10^{-4}$ 1.0
10D 500 RBF 0.6 0.01 $1\times 10^{-4}$ 1.0
64D (Power) 500 RBF 215.0 $\times 10^{-4} 1×1$$\times 10^{-4}$ $U[0.9,1.1]$
Table 12: Transformer Architecture Hyperparameters.
Input Dim Embed Dim Encoder Layers Heads FFN Dim Input Norm
1D 128 1 4 512 Uniform
2D 128 1 4 512 Uniform
5D 64 2 8 1024 Uniform
10D 32 2 8 1024 Uniform
64D 64 4 8 1024 Standard
Table 13: CNN Architecture Hyperparameters.
Input Dim Embed Dim Layers Kernel Size Input Norm
1D 32 1 5 Uniform
2D 32 1 5 Uniform
5D 32 4 5 Uniform
10D 32 4 5 Uniform
64D 32 4 5 Standard
Table 14: Recurrent (RNN/LSTM) Architecture Hyperparameters.
Input Dim Embed Dim Recurrent Layers Attention Heads Dropout
1D 64 1 4 0.1
10D 64 4 8 0.1
Table 15: Optimization Hyperparameters.
Setting Epochs Steps/Epoch Batch Size LR Warmup Epochs
1D (All) 100 500 16 $1\times 10^{-3}$ 25
2D (All) 100 500 16 $1\times 10^{-3}$ 25
5D (All) 200 500 32 $1\times 10^{-3}$ 50
10D (All) 200 500 16 $1\times 10^{-3}$ 50
64D (Power) 200 500 32 $1\times 10^{-3}$ 50
Recurrent models process the set as a sequence and then use an attention mechanism (VA or DVA) to aggregate context.
Optimization and Training
All models were trained using the AdamW optimizer and a cosine-annealing learning rate schedule with linear warmup. The loss function used was the Bar Distribution (Riemannian) Negative Log-Likelihood.
Appendix J Scaling to further dimensions
Figure 15: Validation Loss on 100-Dimensional Data. Performance of the Transformer model utilizing Decoupled Value Attention (DVA) trained over 200 epochs. The $x$-axis represents the cumulative number of training points observed (in billions). The model demonstrates stable convergence on the high-dimensional regression task. The $N_{context}=1500$ and RBF prior was used with a fixed length scale. It took 81.5 hours on NVIDIA 4500ADA GPU having 24GB VRAM.
Appendix K Empirical Comparison with Post-Hoc Localization
To empirically validate our architectural approach against the post-hoc localization method proposed by nagler23a, we conducted a direct comparison. We applied two post-hoc localization strategies to a pre-trained VA PFN at inference time: (1) a k-Nearest Neighbors (k-NN) filter, where only the $k$ closest context points are used, and (2) an Exponential distance filter, where context points are selected based on a distance threshold controlled by a decay factor $\gamma$. The results, shown in Figure 16, provide two key insights.
  • In the low-dimensional (1D) setting, where the VA model is able to learn a useful representation, post-hoc localization is effective in reducing baseline error, and even slightly outperforming DVA with certain neighborhood hyperparameters. This confirms the efficacy of the localization principle in low-dimensional regimes where the base model has learned a meaningful signal.
  • In the high-dimensional (10D) setting, the VA PFN completely fails to learn from the data, with its MSE remaining high and flat regardless of context size. Consequently, applying post-hoc localization to this poorly trained model provides no benefit whatsoever; filtering an uninformative context set is ineffective. In contrast, DVA’s architectural locality enables successful learning from the start, with its error decreasing consistently as more context is provided.
This experiment empirically validates our central hypothesis: post-hoc methods cannot rescue a model that has failed to learn a meaningful spatial representation during training. This underscores the necessity of an architectural solution like DVA for scaling PFNs to high-dimensional regression tasks.
(a) Post-Hoc Localization Performance (1D)
(b) Post-Hoc Localization Performance (10D)
Figure 16: Comparison of DVA with a baseline VA model augmented by post-hoc localization methods. In 1D, post-hoc methods are effective with correct choice of neighborhood parameters. However, in 10D, they fail to improve the poorly trained VA model, while DVA learns successfully.
K.1 Sensitivity Analysis of Post-Hoc k-NN Localization
While post-hoc localization can improve upon a baseline VA model in 1D, its performance is critically dependent on the choice of the hyperparameter $k$ (the number of neighbors). To investigate this dependency, we conducted a simple sensitivity analysis in the 1D setting. We fixed the context size at $N_{\text{context}}=30$ and evaluated the MSE of the VA + Post-Hoc k-NN model for $k\in[1,40]$. We performed this experiment on a RBF prior with length-scale $0.6$.
Figure 17: Sensitivity of Post-Hoc k-NN performance to the choice of $k$ on datasets with different smoothness levels. The optimal $k$ is data-dependent, highlighting the challenge of tuning post-hoc methods. DVA achieves robust, strong performance without requiring such tuning.
The results, shown in Figure 17, reveal a crucial challenge for post-hoc methods. The performance curve for k-NN exhibits a classic U-shape, representing a bias-variance trade-off: small $k$ leads to high variance (overfitting to noisy neighbors), while large $k$ leads to high bias (oversmoothing). Critically, the optimal value of $k$ is data-dependent. This demonstrates that post-hoc localization is not a “set-and-forget” solution. Achieving its optimal performance would require a new, costly hyperparameter search for $k$ for each new function class, undermining the inference-time efficiency of PFNs.